u cos y sinh x, sin y cosh x satisfy laplace equation?



1
Expert's answer
2021-07-16T08:16:28-0400
"u=\cos y\sinh x+\sin y\cosh x"

"u_x=\cos y\cosh x+\sin y\sinh x"

"u_{xx}=\cos y\sinh x+\sin y\cosh x"


"u_y=-\sin y\sinh x+\cos y\cosh x"

"u_{yy}=-\cos y\sinh x-\sin y\cosh x"

Laplace equation


"u_{xx}+u_{yy}=0"

Check


"\cos y\sinh x+\sin y\cosh x"


"+(-\cos y\sinh x-\sin y\cosh x)=0,True"

Therefore the function "u=\cos y\sinh x+\sin y\cosh x" satisfies the Laplace equation.